# Activated LoRA: Fine-tuned LLMs for Intrinsics

### v4

@@ Abstract
Low-Rank Adaptation (LoRA) has emerged as a highly efficient framework for finetuning the weights of large foundation models, and has become the go-to method for data-driven customization of LLMs. Despite the promise of highly customized behaviors and capabilities, switching between relevant LoRAs in a multiturn setting is inefficient, as the key-value (KV) cache of the entire turn history must be recomputed with the LoRA weights before generation can begin. To address this problem, we propose Activated LoRA (aLoRA), an adapter architecture which modifies the LoRA framework to only adapt weights for the tokens in the sequence after the aLoRA is invoked. This change crucially allows aLoRA to accept the base model’s KV cache of the input string, meaning that aLoRA can be instantly activated whenever needed in a chain without recomputing the cache. This enables building what we call intrinsics, i.e. specialized models invoked to perform well-defined operations on portions of an input chain or conversation that otherwise uses the base model by default. We train a set of aLoRA-based intrinsics models, demonstrating competitive accuracy with standard LoRA while achieving significant inference benefits. The codebase is at https://github.com/IBM/activated-lora.111pip install alora
@@ 1 Introduction
The rapid adoption of large language models (LLMs) has catalyzed significant advancements in natural language processing tasks, from text generation to knowledge extraction. However, adapting these models to specific tasks or domains often demands fine-tuning their immense parameter space, a process that is computationally expensive and difficult to scale. Low-Rank Adaptation (LoRA)(Hu et al., 2021) has addressed these challenges by introducing a parameter-efficient method for fine-tuning (Houlsby et al., 2019), enabling highly customized model behavior without the need to retrain or modify the entire model. By optimizing a small subset of low-rank matrices, LoRA has emerged as a widely-used lightweight and effective alternative for task-specific customization, particularly for large foundation models such as LLMs, and popular LoRA finetuning services are offered by both corporate and open-source LLM providers (OpenAI, 2024; TogetherAI, 2024; Predibase, 2024). While large models perhaps have less need of finetuning, small models continue to see strong benefits from LoRA adapters Wang et al. (2025); Danilevsky et al. (2025).
@@
Motivated by this, significant work has been done to enable highly efficient serving of and inference with LoRAs, e.g. Sheng et al. (2023); Brüel-Gabrielsson et al. (2024), with vLLM Kwon et al. (2023) incorporating and further optimizing many of these algorithms in its popular, state-of-the-art inference engine. Yet, while LoRA excels in static or single-task scenarios, in modern applications a wide mix of skills are typically needed, e.g. in multiturn interactions and agentic applications Wang et al. (2024); Yang et al. (2024). Crucially, LoRAs often must be carefully tuned to avoid forgetting/degradation of performance on the wide variety of tasks the base model already performs well on. In applications where dynamic switching between multiple specialized skills would be advantageous (e.g. agentic or reasoning pipelines), this strategy breaks down. Additionally, multi-task LoRA training is significantly more difficult than training a LoRA for a single task, and this approach would be inherently non-modular, requiring retraining to incorporate further abilities later on.
@@
In an ideal world, we would like to be able to seamlessly switch—within the same chat/agentic pipeline—between arbitrary sets of pretrained LoRA configurations for specialized tasks, while keeping the base model as-is for most interactions in the sequence. Unfortunately, the LoRA architecture is not well-suited to this regime, creating major inference inefficiencies. Specifically, for each switch between LoRA adapters, it becomes necessary to recompute the representation of all context tokens prior to generation with the LoRA. In the case of the popular attention mechanism introduced by Vaswani et al. (2017), such a representation takes the form of the key-value (KV) cache of such prior tokens, but more generally, it is a problem that reoccurs even in different architectures; for example, in a state-space model, the state representation of the tokens prior to generation would need to be recomputed if the matrices $A$ and $B$ are updated by a low rank correction. This recalculation introduces significant latency, GPU memory, and computational overhead that all scale with the length of the context that must be prefilled, especially as LLMs increasingly work with documents, files, or reasoning/agentic chains in the many (sometimes hundreds of) thousands of tokens Liu et al. (2025; 2024; 2023) (even millions for software engineering Jiang et al. (2025)). This limits LoRA’s usability in scenarios where rapid transitions between specialized behaviors are essential.
@@
The concept of activated LoRA takes an opinionated view on how such differentiated performance could be attained. As inspiration, we note that instructions in LLMs can appear in many places in a prompt; Figure 1 illustrates an example of two such places: an “early prompting” case where the instruction precedes the content, or a “late prompting” case where the content precedes the instruction. In the latter, the instruction does not have to be revealed ahead of time, making the representation of the content (KV cache) potentially reusable for or from other tasks. We note that these paradigms can take on many shapes; for example, prefix tuning methods such as Li and Liang (2021) explicitly use an early prompting framework, tuning this early prompt. Recently, a set of real-world tasks (and trained adapters) conforming to the late-prompting intrinsics regime was presented in Danilevsky et al. (2025), in the context of building performant and robust RAG Lewis et al. (2020) pipelines. We believe that a vast array of intrinsics-style finetuning tasks can be created with applications throughout the agentic, chat, and reasoning spaces.
Note that a LoRA adapter, like the prefix tuning example above, shares the downsides of the early prompting concept; the LoRA-adapted LLM’s internal representation of the content (KV cache) is fit-for-purpose, specific to the LoRA adapter and not reusable by or from the base model or other adapters.
@@
Addressing this inability to reuse the base model representation of the tokens prior to generation, we introduce Activated LoRA (aLoRA), a novel extension of the LoRA framework designed to only adapt the model’s weights for tokens encountered after activation, allowing base model KV cache for prior context to be reused (Figure 2). By decoupling the adaptation process from the need to recompute the input’s representation, aLoRA facilitates instantaneous switching between specialized models (intrinsics) while maintaining seamless interaction with the base model. This innovation not only reduces computational costs, but also unlocks new possibilities for deploying highly modular, task-specific behaviors within complex workflows.
In our experiments, we demonstrate significant speedups for aLoRA vs LoRA on the state-of-the-art inference engine vLLM Zhu et al. (2023), and show that aLoRA adapters do not lose accuracy versus LoRA on a collection of benchmark finetuning tasks and a collection of real-world “intrinsics” tasks from Danilevsky et al. (2025).
@@
Figure 2: Computation and memory pattern of (a) LoRA vs. (b) aLoRA used as evaluators of an answer given by a base model. (1) prompt is input to the base model, which generates answer, (2) prompt + answer is input to the both intrinsics in parallel, which generate eval_1 and eval_2 respectively. Narrow rectangles denote tokens and wide rectangles denote the KV cache.
@@ 1.1 Background
Modern decoder-only causal LLMs are transformers containing a sequence of decoder layers, each of which typically contain MLP layers that operate on each token’s representation independently, and a causal attention mechanism that allows tokens to attend to representations of prior tokens in the same layer (see Figure 9 in the appendix). We review the ubiquitous softmax attention mechanism introduced by Vaswani et al. (2017) and its connection to the KV cache, as well as LoRA adapters. [^2] Extension of this to other forms of attention is straightforward, but we leave this treatment to future work. Further, note that LoRA and aLoRA adapters can be applied to MLP blocks as well, but since it is typical practice to not adapt the MLP blocks we focus the presentation on the attention blocks only.
@@
LLM inference and the KV cache: LLMs, using a causal attention mask Radford et al. (2019), generate tokens autoregressively one at a time. Suppose we have already generated tokens $11,\dots,t-$ 1 with corresponding hidden-state embeddings $x_{1},\dots,x_{t-1}$. Whenever these tokens were processed (either during prefill or generation) their keys and values were already computed:
@@
Because the entire matrix of past keys $K_{1:t-1}$ and values $V_{1:t-1}$ is precomputed and stored in the “KV cache,” at each new step $t$ we only pay the cost of computing one new row for each of $Q$, $K$, and $V$, and performing a single-row softmax and weighted sum over the remaining cached $K$ and $V$. As shown in Vaswani et al. (2017); Radford et al. (2019), this reduces what would have been an $\mathcal{O}(t^{2}d_{k})$ full-matrix multiply down to only $\mathcal{O}(t\,d_{k})$ operations at each step, yielding massive speedups.
LoRA: LoRA adapts the attention weights $W^{Q}$, $W^{K}$, and $W^{V}$ by replacing them with $W^{Q}+\Delta^{Q}$, $W^{K}+\Delta^{K}$, and $W^{V}+\Delta^{V}$, where $\Delta^{Q},\Delta^{K},\Delta_{V}$ are rank $r$ matrices. This yields
@@
This lowers the number of parameters, making finetuning significantly more efficient (Hu et al., 2021). If the LoRA is active for the entire chain, then the KV cache-based inference applies and generation is efficient. If, on the other hand, any part of the input was generated or prefilled by the base model or another LoRA, then $K$ and $V$ for the LoRA are different than the corresponding $K$ and $V$ for the base model, hence any existing base model KV cache cannot be used by the LoRA, meaning the KV cache for the entire (potentially very long) input must be recomputed. To clarify, this involves passing the input through all layers of the transformer in sequence–not just the adapted attention layers–since modifications to attention blocks modify the inputs to all downstream layers and tokens.
@@ 2 Activated LoRA
Just as in LoRA, our aLoRA architecture adapts the attention weights $W^{Q}$, $W^{K}$, and $W^{V}$ by replacing them with $W^{Q}+\Delta^{Q}$, $W^{K}+\Delta^{K}$, and $W^{V}+\Delta^{V}$, where $\Delta^{Q},\Delta^{K},\Delta^{V}$ are rank $r$ matrices. The difference lies in how these adapted weights are used. We assume that the the default generation model for the chat is the base model, and that intrinsics only operate on these base model generations. As a result, we can assume that the base model has precomputed a KV cache for the input context (the blue region in Figure 1).
@@
For the causal decoder-only transformers we consider, the keys and values (actually, all internal states) prior to $t_{\mathrm{invoke}}$ are identical for the base model and any aLoRA adapter model using (6). Specifically, $K^{base}_{1:t_{\mathrm{invoke}}-1}=K^{adapter}_{1:t_{\mathrm{invoke}}-1}$, and $V^{base}_{1:t_{\mathrm{invoke}}-1}=V^{adapter}_{1:t_{\mathrm{invoke}}-1}$. Inference with the aLoRA adapted model can be done causally one token at a time (by simply increasing $t$ in (6) iteratively) with KV cache reuse.
@@
Consider invoking an adapter with $T_{cache}$ tokens of input for which a base model KV cache exists and $T_{new}\ll T_{cache}$ input tokens without cache. [^3] We presume that $T_{new}$ is the invocation sequence and following. The first token generated by the aLoRA adapter requires $O(T_{cache}T_{new})$ operations, while LoRA requires $O((T_{cache}+T_{new})^{2})$. Furthermore, aLoRA must maintain only $O(T_{new})$ additional KV cache memory, while LoRA requires additional $O(T_{cache}+T_{new})$. If $N$ distinct adapters share the first $T_{cache}$ tokens as input, aLoRA costs become $O(NT_{cache}T_{new})$ and $O(NT_{new})$, while LoRA has $O(N(T_{cache}+T_{new})^{2})$ and $O(N(T_{cache}+T_{new}))$.
As the context length and number of concurrent adapters rise, the computation and memory advantages of aLoRA scale linearly. Note that while methods exist for KV cache compression that can reduce overall memory footprint, e.g. Li et al. (2024), and various methods exist increasing the practical scalability of attention, KV cache reuse is still a major factor in practice as context length increases.
@@
    Multiple adapters reuses the same cache, e.g. when checking base model outputs on multiple axes. This regime creates the most dramatic advantage for aLoRA over LoRA.
In the results, we explore concrete intrinsics which follow these patterns. Note that any KV cache created by aLoRA adapter for their own generations are not reusable by other adapters or the base model, as these tokens come after the adapter weights are turned on. If these strings needs to be input into other adapters or the base model, they should be prefilled again. This is a fairly minor concern, however, as intrinsics generations are often short and per token, prefill is much faster than generation.
Invocation: We found it useful to demarcate the activation of the aLoRA adapter via a short invocation token sequence. Advantages are elaborated on in Appendix A include allowing more space for the adapted weights to process the input. Typically, this sequence will be appended to the prompt prior to generation, just as a “generation prompt” is typically appended to prompts for instruct-tuned model (e.g. <|assistant|> or some other string). While the aLoRA will often be invoked programmatically, this design in principle allows for the base model (or other intrinsics) to call the aLoRA model themselves (we do not explore this behavior in the current work).

### v5

@@ Abstract
Low-Rank Adaptation (LoRA) has emerged as a highly efficient framework for finetuning the weights of large foundation models, and has become the go-to method for data-driven customization of LLMs. Despite the promise of highly customized behaviors and capabilities, switching between relevant LoRAs in a multiturn setting is inefficient, as the key-value (KV) cache of the entire turn history must be recomputed with the LoRA weights before generation can begin. To address this problem, we propose Activated LoRA (aLoRA), an adapter architecture which modifies the LoRA framework to only adapt weights for the tokens in the sequence after the aLoRA is invoked. This change crucially allows aLoRA to accept the base model’s KV cache of the input string, meaning that aLoRA can be instantly activated whenever needed in a chain without recomputing the prior keys and values. This enables building what we call intrinsics, i.e. specialized models invoked to perform well-defined operations on portions of an input chain or conversation that otherwise uses the base model by default. We train a set of aLoRA-based intrinsics models, demonstrating competitive accuracy with standard LoRA while significantly improving inference efficiency. We contributed our Activated LoRA implementation to the Huggingface PEFT library. [^1] https://github.com/huggingface/peft
@@ 1 Introduction
The rapid adoption of large language models (LLMs) has catalyzed significant advancements in natural language processing tasks, from text generation to knowledge extraction. However, adapting these models to specific tasks or domains often demands finetuning their immense parameter space, a process that is computationally expensive and difficult to scale. Low-Rank Adaptation (LoRA)[13] has addressed these challenges by introducing a parameter-efficient method for fine-tuning [12], enabling highly customized model behavior without the need to retrain or modify the entire model. By optimizing a small subset of low-rank matrices, LoRA has emerged as a widely-used lightweight and effective alternative for task-specific customization [25], particularly for large foundation models such as LLMs, and popular LoRA finetuning services are offered by both corporate and open-source LLM providers [26, 36, 28]. While large models perhaps have less need for finetuning, small models continue to see strong benefits from LoRA adapters [40, 5].
@@
Motivated by this, significant work has been done to enable highly efficient serving of and inference with LoRAs, e.g. [34, 2], with vLLM [16] incorporating and further optimizing many of these algorithms in its popular, state-of-the-art inference engine. Yet, while LoRA excels in static or single-task scenarios, in modern applications a wide mix of skills is typically needed, e.g. in multiturn interactions and agentic applications [41, 44]. Crucially, LoRAs often must be carefully tuned to avoid forgetting/degradation of performance on the wide variety of tasks the base model already performs well on. In applications where dynamic switching between multiple specialized skills would be advantageous (e.g. agentic or reasoning pipelines), this strategy breaks down. Additionally, multi-task LoRA training is significantly more difficult than training a LoRA for a single task, and this approach would be inherently non-modular, requiring retraining to incorporate further abilities later on.
@@
In an ideal world, we would like to be able to seamlessly switch—within the same chat/agentic pipeline—between arbitrary sets of pretrained LoRA configurations for specialized tasks, while keeping the base model as-is for most interactions in the sequence. Unfortunately, the LoRA architecture is not well-suited to this regime, creating major inference inefficiencies. Specifically, for each switch between LoRA adapters, it becomes necessary to recompute the representation of all context tokens prior to generation with the LoRA. In the case of the popular attention mechanism introduced by [37], such a representation takes the form of the key-value (KV) cache of such prior tokens [^2] In this work, we use the term “(KV) cache” to denote the set of saved keys and values for prior tokens in the LLM context, noting that our approach is not limited to attention types that use keys and values. We use “(KV) cache reuse” to mean (re)use of stored keys and values from an LLM call (e.g. base model generation) in a subsequent LLM call (e.g. the aLoRA generation) whose context shares a prefix with the original context and/or generation., but more generally, it is a problem that reoccurs even in different architectures; for example, in a state-space model, the state representation of the tokens prior to generation would need to be recomputed if the matrices $A$ and $B$ are updated by a low rank correction. This recalculation introduces significant latency, GPU memory, and computational overhead that all scale with the length of the context that must be prefilled, especially as LLMs increasingly work with documents, files, or reasoning/agentic chains in the many (sometimes hundreds of) thousands of tokens [24, 21, 22] (even millions for software engineering [14]). This limits LoRA’s usability in scenarios where rapid transitions between specialized behaviors are essential.
@@
The concept of activated LoRA takes an opinionated view on how such differentiated performance could be attained. As inspiration, we note that instructions in LLMs can appear in many places in a prompt; Figure 1 illustrates an example of two such places: an “early prompting” case where the instruction precedes the content, or a “late prompting” case where the content precedes the instruction. In the latter, the instruction does not have to be revealed ahead of time, making the representation of the content (KV cache) potentially reusable for or from other tasks. We note that these paradigms can take on many shapes; for example, prefix tuning methods such as [18] explicitly use an early prompting framework, tuning this early prompt. Recently, a set of real-world tasks (and trained adapters) conforming to the late-prompting intrinsics regime was presented in [5], in the context of building performant and robust RAG [17] pipelines. We believe that a vast array of intrinsics-style finetuning tasks can be created with applications throughout the agentic, chat, and reasoning spaces.
A LoRA adapter, like the prefix tuning example above, shares the downsides of the early prompting concept; the LoRA-adapted LLM’s internal representation of the content (KV cache) is fit-for-purpose, specific to the LoRA adapter and not reusable by or from the base model or other adapters. Note that while late-prompt tuning approaches do exist and may be viable in some cases, they tend to underperform prefix tuning (early-prompt tuning), which in turn significantly underperforms LoRA-based training [10]. As a result, a trainable late-prompting style adapter is needed to close this gap.
@@
We therefore introduce Activated LoRA (aLoRA), a novel extension of the LoRA framework designed to only adapt the model’s weights for tokens encountered after activation, allowing base model KV cache for prior context to be reused (Figure 2). By decoupling the adaptation process from the need to recompute the input’s representation, aLoRA facilitates instantaneous switching between specialized models (intrinsics) while maintaining seamless interaction with the base model. This innovation not only reduces computational costs, but also unlocks new possibilities for deploying highly modular, task-specific behaviors within complex workflows.
In our experiments, we demonstrate significant speedups for aLoRA vs LoRA on the state-of-the-art inference engine vLLM [45], and show that aLoRA adapters do not lose accuracy versus LoRA on a collection of benchmark finetuning tasks and a collection of real-world “intrinsics” tasks from [5].
Relationship to other LoRA variants. QLoRA [7] proposes quantizing the weights in the base model while keeping the adapter higher precision in training and inference, which dramatically reduces overall memory costs. This can be directly applied to Activated LoRA as well [^3] Implementation already supported in the Huggingface PEFT library., with KV cache reuse still possible if the base model inference also uses the same or sufficiently similar quantization. Experiments with this are beyond the scope of the present work. DoRA [23] adapters are similar to LoRA with a different decomposed low rank weight matrix, with magnitude and direction of the adaptation parameterized separately. In principle, extension to DoRA-style adapters (“Activated DoRA”) is immediate along the same principles as aLoRA. That said, inference with DoRA is less efficient than LoRA, so it is typically recommended [25] that DoRA adapters be merged into the base model weights — a procedure that is not possible for our “activated” approach due to the selective application of the adapter weights.
@@
Figure 2: Computation and memory pattern of (a) LoRA vs. (b) aLoRA used as evaluators of an answer given by a base model. (1) prompt is input to the base model, which generates answer, (2) prompt + answer is input to both intrinsics in parallel, which generate eval_1 and eval_2 respectively. Narrow rectangles denote tokens and wide rectangles denote the KV cache.
@@ 1.1 Background
Modern decoder-only causal LLMs are transformers containing a sequence of decoder layers, each of which typically contain MLP layers that operate on each token’s representation independently, and a causal attention mechanism that allows tokens to attend to representations of prior tokens in the same layer (see Figure 9 in the appendix). We review the ubiquitous softmax attention mechanism introduced by [37] and its connection to the KV cache, as well as LoRA adapters. [^4] Extension of this to other forms of attention is straightforward, but we leave this treatment to future work. Further, note that LoRA and aLoRA adapters can be applied to MLP blocks as well, but since it is typical practice to not adapt the MLP blocks we focus the presentation on the attention blocks only.
@@
LLM inference and the KV cache: LLMs, using a causal attention mask [29], generate tokens autoregressively one at a time. Suppose we have already generated tokens $1,\dots,t-1$ with corresponding hidden‐state embeddings $x_{1},\dots,x_{t-1}$. Whenever these tokens were processed (either during prefill or generation) their keys and values were already computed:
@@
Because the entire matrix of past keys $K_{1:t-1}$ and values $V_{1:t-1}$ is precomputed and stored in the “KV cache,” at each new step $t$ we only pay the cost of computing one new row for each of $Q$, $K$, and $V$, and performing a single‐row softmax and weighted sum over the remaining cached $K$ and $V$. As shown in [37, 29], this reduces what would have been an $\mathcal{O}(t^{2}d_{k})$ full‐matrix multiply down to only $\mathcal{O}(t\,d_{k})$ operations at each step, yielding massive speedups.
LoRA: LoRA adapts the attention weights $W^{Q}$, $W^{K}$, and $W^{V}$ by replacing them with $W^{Q}+\Delta^{Q}$, $W^{K}+\Delta^{K}$, and $W^{V}+\Delta^{V}$, where $\Delta^{Q},\Delta^{K},\Delta^{V}$ are rank $r$ matrices. This yields
@@
This lowers the number of parameters, making finetuning significantly more efficient [13]. If the LoRA is active for the entire chain, then the KV cache-based inference applies and generation is efficient. If, on the other hand, any part of the input was generated or prefilled by the base model or another LoRA, then $K$ and $V$ for the LoRA are different than the corresponding $K$ and $V$ for the base model, hence any existing base model KV cache cannot be used by the LoRA, meaning the KV cache for the entire (potentially very long) input must be recomputed. To clarify, this involves passing the input through all layers of the transformer in sequence–not just the adapted attention layers–since modifications to attention blocks modify the inputs to all downstream layers and tokens.
@@ 2 Activated LoRA
Just as in LoRA, our aLoRA architecture adapts the attention weights $W^{Q}$, $W^{K}$, and $W^{V}$ by replacing them with $W^{Q}+\Delta^{Q}$, $W^{K}+\Delta^{K}$, and $W^{V}+\Delta^{V}$, where $\Delta^{Q},\Delta^{K},\Delta^{V}$ are rank $r$ matrices. The difference lies in how these adapted weights are used. We assume that the default generation model for the chat is the base model, and that intrinsics only operate on these base model generations. As a result, we can assume that the base model has precomputed a KV cache for the input context (the blue region in Figure 1).
@@
For the causal decoder-only transformers we consider, the keys and values (actually, all internal states) prior to $t_{\mathrm{invoke}}$ are identical for the base model and any aLoRA adapter model using (6). Specifically, $K^{\mathrm{base}}_{1:t_{\mathrm{invoke}}-1}=K^{\mathrm{adapter}}_{1:t_{\mathrm{invoke}}-1}$, and $V^{\mathrm{base}}_{1:t_{\mathrm{invoke}}-1}=V^{\mathrm{adapter}}_{1:t_{\mathrm{invoke}}-1}$. Inference with the aLoRA adapted model can be done causally one token at a time (by simply increasing $t$ in (6) iteratively) with KV cache reuse.
@@
Consider invoking an adapter with $T_{\mathrm{cache}}$ tokens of input for which a base model KV cache exists and $T_{\mathrm{new}}\ll T_{\mathrm{cache}}$ input tokens without cache. [^5] We presume that $T_{\mathrm{new}}$ is the invocation sequence and following. The first token generated by the aLoRA adapter requires $O(T_{\mathrm{cache}}T_{\mathrm{new}})$ operations, while LoRA requires $O((T_{\mathrm{cache}}+T_{\mathrm{new}})^{2})$. Furthermore, aLoRA must maintain only $O(T_{\mathrm{new}})$ additional KV cache memory, while LoRA requires additional $O(T_{\mathrm{cache}}+T_{\mathrm{new}})$. If $N$ distinct adapters share the first $T_{\mathrm{cache}}$ tokens as input, aLoRA costs become $O(NT_{\mathrm{cache}}T_{\mathrm{new}})$ and $O(NT_{\mathrm{new}})$, while LoRA has $O(N(T_{\mathrm{cache}}+T_{\mathrm{new}})^{2})$ and $O(N(T_{\mathrm{cache}}+T_{\mathrm{new}}))$.
As the context length and number of concurrent adapters rise, the computation and memory advantages of aLoRA scale linearly. Note that while methods exist for KV cache compression that can reduce overall memory footprint, e.g. [19], and various methods exist increasing the practical scalability of attention, KV cache reuse is still a major factor in practice as context length increases.
@@
    Multiple adapters reuse the same cache, e.g. when checking base model outputs on multiple axes. This regime creates the most dramatic advantage for aLoRA over LoRA.
In the results, we explore concrete intrinsics which follow these patterns. Note that any KV cache created by aLoRA adapter for their own generations are not reusable by other adapters or the base model, as these tokens come after the adapter weights are turned on. If these strings needs to be input into other adapters or the base model, they should be prefilled again. This is a fairly minor concern, however, as intrinsics generations are often short, and per token, prefill is much faster than generation.
Invocation: We found it useful to demarcate the activation of the aLoRA adapter via a short invocation token sequence. Advantages are elaborated on in Appendix A and include allowing more space for the adapted weights to process the input. Typically, this sequence will be appended to the prompt prior to generation, just as a “generation prompt” is typically appended to prompts for instruct-tuned model (e.g. <|assistant|> or some other string). While the aLoRA will often be invoked programmatically, this design in principle allows for the base model (or other intrinsics) to call the aLoRA model themselves (we do not explore this behavior in the current work).